# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

### v1

@@
Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in
[…]
${}^{1}\,$ Department of Physics, Indian Institute of Technology Hyderabad, Kandi, Sangareddy 502285, Telangana, India
The singlet-doublet vector-like fermion dark matter model has been extensively studied in the literature over the past decade. An important parameter in this model is the singlet-doublet mixing angle ($\sin\theta$). All the previous studies have primarily focused on annihilation and co-annihilation processes for obtaining the correct dark matter relic density, assuming that the singlet and doublet components decouple at the same epoch. In this work, we demonstrate that this assumption holds only for larger mixing angles with a dependency on the mass of the dark matter. However, it badly fails for the mixing angle $\sin\theta<0.05$. We present a systematic study of the parameter space of the singlet-doublet dark matter relic, incorporating annihilation, co-annihilation, and, for the first time, co-scattering processes. Additionally, the freeze-in parameter space is also explored. We found that due to the inclusion of co-scattering processes, the correct relic density parameter space is shifted towards the detection sensitivity range of the LHC and MATHUSLA via displaced vertex signatures.
@@ 1 Introduction
A simplest extension of the SM for a DM candidate is a vector-like singlet fermion: $\chi$ with an additional $\mathcal{Z}_{2}$ symmetry under which $\chi$ is odd while all other particles are even [8]. It has only a dimension five effective coupling with the SM Higgs boson , i.e. $(\bar{\chi}\chi H^{\dagger}H)/\Lambda$, where $\Lambda$ is the cutoff scale. A very large coupling is required to satisfy the correct relic density. However, the direct detection experiments exclude such large couplings. In the singlet fermionic DM model, only DM mass at the Higgs resonance ($M_{\rm DM}=M_{H}/2$) survives the direct detection constraints. On the other hand, if the SM is extended with a $\mathcal{Z}_{2}$-odd vector-like fermion doublet, then the neutral component of the doublet can be a viable DM candidate[8]. However, due to large gauge coupling, the annihilation cross-section is so large that the DM becomes under-abundant in the low mass region. The relic is satisfied at a very high DM mass, typically at the TeV scale. Moreover, the doublet DM model is ruled out by the current direct detection constraints [6]. This problem can be resolved if the singlet is combined with the vector-like doublet fermion, which are odd under a $\mathcal{Z}_{2}$ symmetry. In this case, the neutral component of the doublet mixes with the singlet to give a singlet-doublet DM (SDDM). By choosing the SD mixing appropriately, one can get the correct relic density in a large range of parameter space, which also satisfies the direct detection constraints. The SDDM has been studied extensively in the literature [7, 9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. An interesting feature of this model is that it introduces co-annihilation processes along with the annihilation among the dark sector particles to give rise the correct relic density of DM. The two most important parameters in the SDDM model are the SD mixing angle ($\sin\theta$) and the mass splitting ($\Delta{M}$) between the DM (dominantly singlet) and the NLSP (dominantly the doublet). For a larger $\sin\theta$ and $\Delta M$, the annihilation process dominantly decides the relic density of the DM. On the other hand, in the small mass-splitting regime, the co-annihilation process becomes dominant. In previous studies of the SDDM [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18], it was generally assumed that the dark sector particles decouple from the thermal bath at the same epoch, independent of $\sin\theta$. However, this assumption may not hold true for small mixing angles; typically, for $\sin\theta\lesssim 0.05$, the DM particles are not in chemical equilibrium with the thermal bath. In particular, the singlet component of the dark sector decouples from the thermal bath at an earlier time, while the doublet component remains in equilibrium for a longer duration due to its interactions with the SM gauge bosons. This implies that for $\sin\theta\lesssim 0.05$, the relic density parameter space estimated in earlier studies is not valid, while they remain valid for $\sin\theta\gtrsim 0.05$
@@
In this paper, we revisit the SDDM relic density in the small mixing limits, typically $10^{-6}\lesssim\sin\theta\lesssim 0.05$. In this range of SD mixing, the singlet can be brought to equilibrium quickly. However, the DM (dominantly singlet) decouples early with a larger abundance. We demonstrate that by incorporating co-scattering or conversion-driven processes [32, 33], the DM relic density can be brought into agreement with the observed value. In particular, the kinetic equilibrium between DM and the thermal bath allows for upscattering processes, where DM particles can be scattered into a higher-mass state within the dark sector, such as the doublet, thereby reducing the singlet’s abundance. Recent studies have explored co-scattering processes in various BSM contexts; see, for instance, [34, 35, 36].
@@
We further explore the DM parameter space by reducing the SD mixing angle. We saw that for $\sin\theta<\mathcal{O}(10^{-6})$, the singlet never reaches equilibrium, and hence the DM relic density can not be realized via the freeze-out mechanism. In such a small mixing angle regime, the DM can be produced non-thermally from the decay of the doublet partner. In Section 3.3, we saw that for DM mass ranging from 1 to 1000 GeV, the required $\sin\theta$ for correct relic density varies from $10^{-13}$ to $10^{-7}$.
@@
We demonstrate that in the small mixing limit, the SDDM can yield distinctive signatures at colliders. The doublet fermion can be produced copiously due to their gauge interaction. Once produced, they can decay to DM and leptons via the singlet doublet mixing, which can give displaced vertex signatures in the LHC [37] and, in future, MATHUSLA [38].
@@
The paper is organized as follows. We discuss the SDDM model in Section 2. In Section 3.1, we study the correct relic parameter space considering annihilation and co-annihilation only. We study the co-scattering effects in Section 3.2. The freeze-in production of DM is discussed in Section 3.3. The constraint on the DM parameter space using the null detection of the DM at direct search experiments is discussed in Section 4. Section 5 discusses the displaced vertex signatures. We finally conclude in Section 6.
@@ 2 The singlet-doublet dark matter model
In this case, the SM is extended with a fermion doublet $\Psi=(\psi^{0}~{}~{}\psi^{-})^{T}$ and a singlet fermion $\chi$, which are stabilized by a discrete symmetry $\mathcal{Z}_{2}$, under which they are odd, and all other SM particles are charged even [8].
@@
$\mathcal{L}$ $\supset$ $i\overline{\Psi}\gamma^{\mu}D_{\mu}\Psi+i\overline{\chi}\gamma^{\mu}\partial_{\mu}\chi-M_{\Psi}\overline{\Psi}\Psi-M_{\chi}\overline{\chi}\chi-y\overline{\Psi}\tilde{H}\chi+h.c.,$ (1)
@@
$V_{\rm scalar}$ $=$ $-\mu^{2}_{H}(H^{\dagger}H)+\lambda_{H}(H^{\dagger}H)^{2}.$ (2)
@@
$H=\begin{pmatrix}0\\ \frac{v+h}{\sqrt{2}}\end{pmatrix}.$ (3)
@@
$\left(\begin{matrix}\psi^{0}\\ \chi\end{matrix}\right)=\left(\begin{matrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{matrix}\right)\left(\begin{matrix}\chi_{0}\\ \chi_{1}\end{matrix}\right),$ (4)
@@
$\tan{2\theta}=\frac{\sqrt{2}yv}{M_{\Psi}-M_{\chi}}.$ (5)
@@
$M_{\chi_{0}}$ $=$ $M_{\Psi}\cos^{2}\theta+\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\sin^{2}\theta,~{}~{}$
$M_{\chi_{1}}$ $=$ $M_{\Psi}\sin^{2}\theta-\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\cos^{2}\theta\equiv M_{\rm DM}$ (6)
with a mass-splitting $\Delta M=M_{\chi_{0}}-M_{\chi_{1}}$, where, $\chi_{1}$ is dominantly the singlet fermion and $\chi_{0}$ is the $\psi^{0}$. From Eq. (1), the interaction among the mass eigenstates ($\chi_{0}$, $\psi^{\pm}$ and $\chi_{1}$), can be expressed as,
$\mathcal{L}_{int}$ $=$ $\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\left[\sin^{2}\theta\overline{\chi_{1}}\gamma^{\mu}Z_{\mu}\chi_{1}+\cos^{2}\theta\overline{\chi_{0}}\gamma^{\mu}Z_{\mu}\chi_{0}\right.$ (7)
@@
$-\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\cos 2\theta_{W}{\psi^{+}}\gamma^{\mu}Z_{\mu}\chi^{-}-e_{0}{\psi^{+}}\gamma^{\mu}A_{\mu}\psi^{-}$
@@
where $e_{0}=0.313$ is the electromagnetic coupling constant, and $\sin\theta_{W}$ is the Weinberg angle.
@@ 3 Dark matter phenomenology
In this Section, we systematically explore the relic density parameter space for SDDM, considering annihilation, co-annihilation, co-scattering, and freeze-in, depending on the $\sin\theta$ and $\Delta{M}$. We divide the parameter space into three regimes: (i) thermal DM relic considering annihilation and co-annihilation only, (ii) thermal DM relic via co-annihilation and co-scattering processes, and (iii) freeze-in production of the DM.
3.1 Thermal dark matter relic considering annihilation and co-annihilation only
@@
${}\frac{dn}{dt}+3\mathcal{H}n=-\langle\sigma v\rangle_{\rm eff}(n^{2}-(n^{\rm eq})^{2})$ (8)
where $n=\sum_{i}n_{i}$ represents the total number density of all the dark sector particles, $n^{\rm eq}$ is the equilibrium number density, and the Hubble parameter is expressed as $\mathcal{H}=\sqrt{\frac{4\pi^{2}g_{*\rho}}{45M_{Pl}}}T^{2}$ with the Planck mass, $M_{pl}=1.22\times 10^{19}$ GeV, and $g_{*\rho}$ is the total relativistic d.o.f contributing to the energy density. $\langle\sigma v\rangle_{\rm eff}$ represents the effective cross-section which takes into account all number changing processes for DM freeze-out is given by
$\langle\sigma v\rangle_{\rm eff}$ $=$ $\frac{g^{2}_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{1}}\chi_{{}_{1}}}+\frac{g_{0}g_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{0}}\chi_{{}_{1}}}(1+\Delta_{\chi_{{}_{0}}})^{{3}/{2}}\exp(-x\Delta_{\chi_{{}_{0}}})$ (9)
@@
where $g_{0},g_{1}$ and $g_{2}$ represent the internal d.o.f. of $\chi_{0},\chi_{1}$ and $\psi^{-}$ respectively and $\Delta_{i}$ stands for the ratio $(M_{i}-M_{\chi_{1}})/M_{\chi_{1}}$ with $M_{i}$ denoting the mass of $\chi_{0}$ and $\psi^{-}$. Here $g_{\rm eff}$ is the effective degree of freedom which is given by:
$g_{\rm eff}=g_{1}+g_{0}(1+\Delta_{\chi_{{}_{0}}})^{3/2}\exp(-x\Delta_{\chi_{{}_{0}}})+g_{2}(1+\Delta_{\psi^{-}})^{3/2}\exp(-x\Delta_{\psi^{-}})$ (10)
and $x$ is the dimensionless parameter $M_{\rm\chi_{1}}/T$. The relic density of DM $\chi_{1}$ can then be obtained as
[…]
@@
Upon analyzing Eq. (9), the first three terms correspond to the annihilation and co-annihilation of $\chi_{1}$, while the last three terms describe the annihilation and co-annihilation processes involving $\chi_{0}$ and $\psi^{-}$. The latter terms are associated with gauge-mediated interactions and are independent of the mixing angle: $\sin\theta$. Consequently, these processes dominate over the first three terms in the regime of smaller values of $\sin\theta$. Therefore, it is essential to verify whether the DM remains in chemical equilibrium with the $\mathcal{Z}_{2}$-odd doublet fermion in this low $\sin\theta$ region. We define a new parameter, $\rho_{1}$, to validate the assumption of chemical equilibrium among the dark sector particles,
$\rho_{1}=\frac{\Omega_{1s}h^{2}}{\Omega_{2s}h^{2}(\rm no~{}coscattering)},$ (12)
where $\Omega_{1s}h^{2}$ is the relic density obtained by solving Eq (8), and $\Omega_{2s}h^{2}$(no co-scattering) is calculated by solving Eqs (LABEL:eq:Y1) and (LABEL:eq:Y2) while excluding co-scattering processes [^1] We use micrOMEGAs 6.1.15[40, 35] to solve the Boltzmann equations, where darkOmega function is used to estimate $\Omega_{1s}h^{2}$ and darkOmegaN is used to estimate $\Omega_{2s}h^{2}$(no co-scattering) while excluding the co-scattering processes.. In Eq. (12), $\rho_{1}=1$ implies all the dark sector particles decouple at the same epoch. In other words, solving Eq. (8) or Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) are equivalent. $\rho_{1}<1$ implies the dark sector particles are decoupling at different epochs. When $\sin\theta$ is small, the singlet decouples earlier than the doublet and produces a relatively larger relic. Therefore, one expects $\rho_{1}\leq 1$.
Figure 1: Allowed parameter space of DM for which DM can be considered to be in equilibrium with SM bath and the $\mathcal{Z}_{2}$-odd doublet. All the points in the plots satisfy the required relic density. The grey-shaded region in the plot shows the LEP bound [41] on the heavy fermion doublet.
We have presented the parameter space in Fig. 1. In this plot, all the points satisfy the correct relic density of DM, and we have used $0.99<\rho_{1}\leq 1$ to assess the validity and domain of applicability of the Eq. (8). In Fig. 1 (left), we have shown all the points in the plane of $\sin\theta$ and $M_{\rm DM}$, and in the right plot, we projected the same points in the plane of $\Delta{M}$ and $M_{\rm DM}$. DM mass is varied in the range of $1-1000$ GeV, $\Delta{M}$ is varied in the range of $1-1000$ GeV, and $\sin\theta$ is varied in the range of ($0.5-10^{-4}$). From Fig. 1 (left), we observe that the value of $\sin\theta$ can be as small as $\sim 5\times 10^{-4}$ for $M_{\rm DM}\sim 250$ GeV. Note that all these points in Fig. 1 correspond to the simultaneous decoupling of singlet and doublet from the thermal bath. As shown in Fig. 1 (right), we exactly reproduced all the points that satisfy correct relic density in the plane of $\Delta{M}$ vs $M_{\rm DM}$ [8].
3.2 Thermal dark matter relic via co-annihilation and
co-scattering processes
@@ 3.2 Thermal dark matter relic via co-annihilation and
We have already demonstrated in Fig. 1 (left) the parameter space for the correct relic of DM in $\sin\theta$-$M_{\rm DM}$ plane, where the singlet and the doublet freeze-out at the same epoch through annihilation and co-annihilation processes. Beyond this regime, the singlet may reach equilibrium. However, in this case, the singlet and doublet will decouple at different epochs. Due to small mixing, the singlet will decouple early, while the doublet components will remain in equilibrium for a longer period due to their gauge interactions. Even though the singlet decouples early, the freeze-out abundance of DM may alter because of the presence of additional processes, such as co-scattering or conversion-driven freeze-out [32, 33], which were subdominant for larger SD mixing. As a result, different sets of evolution equations for the dark sector particles have to be solved to get the relic density. In the SDDM model, we define two dark sectors: (a) sector 1, which contains $\chi_{1}$, and (b) sector 2, which contains $\chi_{0},\psi^{-}$, whereas all the SM particles are assigned as sector 0. The singlet fermion, being the lightest among the dark sector particles, serves as the DM candidate. By defining the comoving number density of sector 1 and sector 2 particles as $Y_{1}$ and $Y_{2}$  [^2] Processes like $\psi^{-}SM\rightarrow\chi_{0}SM$ will keep the neutral and charged components of the doublet in equilibrium always which makes the abundances of sector 2 particles to be the same., respectively, the general coupled BEs for their evolution are given as
$\frac{dY_{1}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{1100}v\rangle(Y_{1}^{2}-{Y_{1}^{eq}}^{2})+\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})\right.$

### v2

@@
Abstract
The singlet-doublet vector-like fermion dark matter model has been extensively studied in the literature over the past decade. An important parameter in this model is the singlet-doublet mixing angle ($\sin\theta$). All the previous studies have primarily focused on annihilation and co-annihilation processes for obtaining the correct dark matter relic density, assuming that the singlet and doublet components decouple at the same epoch. In this work, we demonstrate that this assumption holds only for larger mixing angles with a dependency on the mass of the dark matter. However, it badly fails for the mixing angle $\sin\theta<0.05$. We present a systematic study of the parameter space of the singlet-doublet dark matter relic, incorporating annihilation, co-annihilation, and, for the first time, co-scattering processes. Additionally, non-thermal productions via the freeze-in and SuperWIMP mechanism are also explored. We found that due to the inclusion of co-scattering processes, the correct relic density parameter space is shifted towards the detection sensitivity range of the LHC and MATHUSLA via displaced vertex signatures.
@@ 1 Introduction
A simplest extension of the SM for a DM candidate is a vector-like singlet fermion: $\chi$ with an additional $\mathcal{Z}_{2}$ symmetry under which $\chi$ is odd while all other particles are even [6]. It has only a dimension five effective coupling with the SM Higgs boson , i.e. $(\bar{\chi}\chi H^{\dagger}H)/\Lambda$, where $\Lambda$ is the cutoff scale. A very large coupling is required to satisfy the correct relic density. However, the direct detection experiments exclude such large couplings. In the singlet fermionic DM model, only DM mass at the Higgs resonance ($M_{\rm DM}=M_{H}/2$) survives the direct detection constraints. On the other hand, if the SM is extended with a $\mathcal{Z}_{2}$-odd vector-like fermion doublet, then the neutral component of the doublet can be a viable DM candidate [6]. However, due to large gauge coupling, the annihilation cross-section is so large that the DM becomes under-abundant in the low mass region. The relic is satisfied at a very high DM mass, typically at the TeV scale. Moreover, the doublet DM model is ruled out by the current direct detection constraints [7]. This problem can be resolved if the singlet is combined with the vector-like doublet fermion, which are odd under a $\mathcal{Z}_{2}$ symmetry. In this case, the neutral component of the doublet mixes with the singlet to give a singlet-doublet DM (SDDM). By choosing the SD mixing appropriately, one can get the correct relic density in a large range of parameter space, which also satisfies the direct detection constraints. The SDDM has been studied extensively in the literature [8, 9, 6, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. An interesting feature of this model is that it introduces co-annihilation processes along with the annihilation among the dark sector particles to give rise the correct relic density of DM. The two most important parameters in the SDDM model are the SD mixing angle ($\sin\theta$) and the mass splitting ($\Delta{M}$) between the DM (dominantly singlet) and the NLSP (dominantly the doublet). For a larger $\sin\theta$ and $\Delta M$, the annihilation process dominantly decides the relic density of the DM. On the other hand, in the small mass-splitting regime, the co-annihilation process becomes dominant. In previous studies of the SDDM [9, 6, 10, 11, 12, 13, 14, 15, 16, 17, 18], it was generally assumed that the dark sector particles decouple from the thermal bath at the same epoch. However, this assumption may not hold true for small mixing angles; typically, for $\sin\theta\lesssim 0.05$, the DM particles are not in chemical equilibrium with the thermal bath. In particular, the singlet component of the dark sector decouples from the thermal bath at an earlier time, while the doublet component remains in equilibrium for a longer duration due to its interactions with the SM gauge bosons. This implies that for $\sin\theta\lesssim 0.05$, the relic density parameter space estimated in earlier studies is not valid, while they remain valid for $\sin\theta\gtrsim 0.05$.
@@
In this paper, we revisit the SDDM relic density in the small mixing limits, typically $\mathcal{O}(10^{-7})\lesssim\sin\theta\lesssim 0.05$  [^1] If the SD mixing is large, typically $\mathcal{O}(1)$ and $\Delta{M}$ ¡ 1 GeV, then the non-perturbative effects would be important in the relic density calculation, see e.g. [37, 38, 39]. Here, for simplicity, we took $\Delta{M}\geq 1$ GeV.. In this range of SD mixing, the singlet can be brought to equilibrium quickly. However, the DM (dominantly singlet) decouples early with a larger abundance. We demonstrate that by incorporating conversion-driven processes [40, 41], the DM relic density can be brought into agreement with the observed value. In particular, the kinetic equilibrium between DM and the thermal bath allows for upscattering processes, where DM particles can be scattered into a higher-mass state within the dark sector, such as the doublet, thereby reducing the singlet’s abundance. Recent studies have explored the conversion-driven processes in various BSM contexts; see, for instance, [42, 43, 44, 45, 46].
@@
We further explore the DM parameter space by reducing the SD mixing angle. We saw that for $\sin\theta<\mathcal{O}(10^{-7})$, the singlet never reaches equilibrium, and hence the DM relic density can not be realized via the freeze-out mechanism. In such a small mixing angle regime, the DM can be produced non-thermally from the decay of the doublet partner. In Section 3.2, we show that for DM mass ranging from 1 to 1000 GeV and $\Delta M\geq 1$ GeV, the required $\sin\theta$ for correct relic density varies from $\mathcal{O}(10^{-11})$ to $\mathcal{O}(10^{-13})$. On the hand, for SD mixing in the range, $\mathcal{O}(10^{-11})<\sin\theta<\mathcal{O}(10^{-7})$, the DM is over produced. For $\sin\theta<\mathcal{O}(10^{-14})$, DM relic can be achieved via the SuperWIMP mechanism [47, 48, 15] from the out-of-equilibrium decay of the doublet components.
@@
In Section 5, we also demonstrate that the charged partner of the SDDM can yield distinctive signatures at colliders. The doublet fermion can be produced copiously due to their gauge interaction. Once produced, they can decay to DM and leptons via the singlet doublet mixing, which can give displaced vertex signatures in the LHC [49] and, in future, MATHUSLA [50].
@@
The paper is organized as follows. We discuss the SDDM model in section 2. In section 3.1.1, we study the correct relic parameter space considering annihilation and co-annihilation only. We study the effects of conversion-driven processes in section 3.1.2. The freeze-in production of DM is discussed in section 3.2. The production of DM relic via SuperWIMP mechanism is discussed in section 3.3. The constraint on the DM parameter space using the null detection of the DM at direct search experiments is discussed in section 4. Section 5 discusses the displaced vertex signatures. We finally conclude in section 6.
@@ 2 The singlet-doublet dark matter model
In this case, the SM is extended with a fermion doublet $\Psi=(\psi^{0}\leavevmode\nobreak\ \leavevmode\nobreak\ \psi^{-})^{T}$ and a singlet fermion $\chi$, which are stabilized by a discrete symmetry $\mathcal{Z}_{2}$, under which they are odd, and all other SM particles are charged even [6].
@@
$\mathcal{L}$ $\supset$ $i\overline{\Psi}\gamma^{\mu}D_{\mu}\Psi+i\overline{\chi}\gamma^{\mu}\partial_{\mu}\chi-M_{\Psi}\overline{\Psi}\Psi-M_{\chi}\overline{\chi}\chi-y\overline{\Psi}\tilde{H}\chi+h.c.,$ (2.1)
@@
$V_{\rm scalar}$ $=$ $-\mu^{2}_{H}(H^{\dagger}H)+\lambda_{H}(H^{\dagger}H)^{2}.$ (2.2)
@@
$H=\begin{pmatrix}0\\ \frac{v+h}{\sqrt{2}}\end{pmatrix}$. (2.3)
@@
$\left(\begin{matrix}\psi^{0}\\ \chi\end{matrix}\right)=\left(\begin{matrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{matrix}\right)\left(\begin{matrix}\chi_{0}\\ \chi_{1}\end{matrix}\right),$ (2.4)
@@
$\tan{2\theta}=\frac{\sqrt{2}yv}{M_{\Psi}-M_{\chi}}.$ (2.5)
@@
$M_{\chi_{0}}$ $=$ $M_{\Psi}\cos^{2}\theta+\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\sin^{2}\theta,\leavevmode\nobreak\ \leavevmode\nobreak\$
$M_{\chi_{1}}$ $=$ $M_{\Psi}\sin^{2}\theta-\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\cos^{2}\theta\equiv M_{\rm DM}$ (2.6)
with a mass-splitting $\Delta M=M_{\chi_{0}}-M_{\chi_{1}}$, where, $\chi_{1}$ is dominantly the singlet fermion and $\chi_{0}$ is the $\psi^{0}$. From Eq. (2.1), the interaction among the mass eigenstates ($\chi_{0}$, $\psi^{\pm}$ and $\chi_{1}$), can be expressed as,
$\mathcal{L}_{int}$ $=$ $\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\left[\sin^{2}\theta\overline{\chi_{1}}\gamma^{\mu}Z_{\mu}\chi_{1}+\cos^{2}\theta\overline{\chi_{0}}\gamma^{\mu}Z_{\mu}\chi_{0}\right$. (2.7)
@@
$-\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\cos 2\theta_{W}{\psi^{+}}\gamma^{\mu}Z_{\mu}\psi^{-}-e_{0}{\psi^{+}}\gamma^{\mu}A_{\mu}\psi^{-}$
@@
where $e_{0}=0.313$ is the electromagnetic coupling constant, and $\theta_{W}$ is the Weinberg angle. Note that in this model, there are three parameters: $M_{\rm DM},\sin\theta,\Delta M$, which decide the relic. In the rest of the draft, we will find the appropriate parameter space for which the correct relic density will be obtained.
@@ 3 Dark matter phenomenology
In this section, we systematically explore the relic density parameter space for SDDM, considering annihilation, co-annihilation, co-scattering, and freeze-in, depending on the $\sin\theta$ and $\Delta{M}$. We divide the parameter space into four regimes: (i) thermal DM relic considering annihilation and co-annihilation only, (ii) thermal DM relic via conversion-driven processes processes, (iii) dark matter production via freeze-in mechanism, and (iv) dark matter production via SuperWIMP mechanism.
3.1 Thermal production of dark matter
@@
${}\frac{dn}{dt}+3\mathcal{H}n=-\langle\sigma v\rangle_{\rm eff}(n^{2}-(n^{\rm eq})^{2})$ (3.1)
where $n=\sum_{i}n_{i}$ represents the total number density of all dark sector particles, $n^{\rm eq}$ is the corresponding equilibrium number density, and the Hubble parameter is expressed as $\mathcal{H}=\sqrt{\frac{4\pi^{3}g_{*\rho}}{45M^{2}_{\rm Pl}}}T^{2}$ with the Planck mass, $M_{\rm Pl}=1.22\times 10^{19}$ GeV, and $g_{*\rho}$ is the total relativistic contribution to the energy density. $\langle\sigma v\rangle_{\rm eff}$ represents the effective cross-section which takes into account all number changing processes for DM freeze-out and is given by
$\langle\sigma v\rangle_{\rm eff}$ $=$ $\frac{g^{2}_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{1}}\chi_{{}_{1}}}+\frac{g_{0}g_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{0}}\chi_{{}_{1}}}(1+\Delta_{\chi_{{}_{0}}})^{{3}/{2}}\exp(-x\Delta_{\chi_{{}_{0}}})$ (3.2)
@@
where $g_{0},g_{1}$ and $g_{2}$ represent the internal d.o.f. of $\chi_{0},\chi_{1}$ and $\psi^{-}$ respectively and $\Delta_{i}$ stands for the ratio $(M_{i}-M_{\chi_{1}})/M_{\chi_{1}}$ with $M_{i}$ denoting the mass of $\chi_{0}$ and $\psi^{-}$. Here $g_{\rm eff}$ is the effective degree of freedom which is given by
$g_{\rm eff}=g_{1}+g_{0}(1+\Delta_{\chi_{{}_{0}}})^{3/2}\exp(-x\Delta_{\chi_{{}_{0}}})+g_{2}(1+\Delta_{\psi^{-}})^{3/2}\exp(-x\Delta_{\psi^{-}})$ (3.3)
and $x$ is the dimensionless parameter defined as $x=M_{\rm\chi_{1}}/T$. The relic density of DM $\chi_{1}$ can then be obtained as
$\Omega_{\chi_{1}}h^{2}\equiv\Omega_{\rm DM}h^{2}=\frac{1.09\times 10^{9}{\rm GeV}^{-1}}{\sqrt{g_{*}}M_{\rm Pl}}\left[\int_{x_{f}}^{\infty}dx\leavevmode\nobreak\ \frac{\langle\sigma v\rangle_{\rm eff}}{x^{2}}\right]^{-1}.$ (3.4)
@@
Upon analyzing Eq. (3.2), the first three terms correspond to the annihilation and co-annihilation of $\chi_{1}$, while the last three terms describe the annihilation and co-annihilation processes involving $\chi_{0}$ and $\psi^{-}$ (The respective processes are listed in Appendices B.1, B.3 and B.2.). The latter terms are associated with gauge-mediated interactions and are independent of the mixing angle: $\sin\theta$. Consequently, these processes dominate over the first three terms in the regime of smaller values of $\sin\theta$. Therefore, it is essential to verify whether the DM remains in chemical equilibrium with the $\mathcal{Z}_{2}$-odd doublet fermion in this low $\sin\theta$ region. Previous studies [6] have implicitly assumed that the dark sector particles ($\chi_{1},\chi_{0},\psi^{-}$) decouple at the same epoch for $\sin\theta\sim\mathcal{O}(10^{-2})$. In this work, we first verify the validity of this assumption and find the lowest value of $\sin\theta$ for which this assumption is valid. In other words, we find for what values of $\sin\theta$ the dark sector particles chemically decouple at the same epoch. We then show that even for small mixing angles, when the dark sector particles chemically decouple at different epochs, the correct relic abundance can still be achieved through processes known as co-scattering. In this mechanism, the chemically decoupled DM continues to scatter with the SM bath, producing heavier dark-sector states. In our scenario, the singlet DM can scatter off SM particles to produce doublet states. A detailed list of such processes is provided in the Appendix B.4. However, such processes are not included in Eq. (3.1).
@@
In order to incorporate such co-scattering processes, we define two dark sectors: (a) sector 1, which contains $\chi_{1}$, and (b) sector 2, which contains $\chi_{0},\psi^{-}$, whereas all the SM particles are assigned as sector 0. The singlet fermion, being the lightest among the dark sector particles, serves as the DM candidate. By defining the comoving number density of sector 1 and sector 2 particles as $Y_{1}\left(=\frac{n_{\chi_{1}}}{s}\right)$ and $Y_{2}\left(=\frac{n_{\chi_{0}}+n_{\psi^{\pm}}}{s}\right)$  [^2] Processes like $\psi^{-}SM\rightarrow\chi_{0}SM$ will keep the neutral and charged components of the doublet in equilibrium always which makes the abundances of sector 2 particles to be the same., respectively, the general coupled BEs for their evolution are given as
$\frac{dY_{1}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{1100}v\rangle(Y_{1}^{2}-{Y_{1}^{\rm eq}}^{2})+\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{\rm eq}}^{2}}{{Y_{2}^{\rm eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{\rm eq}Y_{2}^{\rm eq})\right$.